OpenStudy (anonymous):

-3( x+1) =x-5+2(5x+5)

OpenStudy (anonymous):

-2??

OpenStudy (anonymous):

the volume of 100 drops of a liquid is 0.1 fl. oz. what is the volume of 10,000 drops?

OpenStudy (anonymous):

What did you get after distributing the -3 and the 2 from the first equation?

OpenStudy (anonymous):

-3x-3=x-5+10x+10 then + 3x to both sides of the equation +3x +3x -3=x-5+13x+20 -3=14x-5+20 -3= 14x+15 -15 -15 -28=14x x=-2

OpenStudy (anonymous):

hi pol

OpenStudy (anonymous):

im going to elaborate , when you get the chance, on the physics question

OpenStudy (anonymous):

-3( x+1) =x-5+2(5x+5) -3x-3=x-5+10x +10 -3x-3=11x+5 -3-5=11x+3x -8=14x x=-8/14 x=-4/7

OpenStudy (anonymous):

Yeah, somehow your +10 became a +20 after you added 3x.

OpenStudy (anonymous):

ok thank you guys so muchhh..too late i had already put in the answer before i asked u guys. dang it.

OpenStudy (anonymous):

i think you would distrubute the -3 to the (x+1) and you would get -3x + -2, and then you would go to the next set of parentheses and distrubute the 2 to the (5x + 5) and you would get 10 x + 10, then you would have -3x + -2 = x -5 + 10x +10, and you would need to move all the (x) values to the left side of the equation, ok well just rember the order operations, and that should help.